a hiker hikes up a mountain 1 mi/h slower than she hikes down the mountain. if it takes her 1.5 hours to hike up the mountain, and only 1 hour to hike down it, how fast does she movein each direction?
Answer by mananth(16946)

a hiker hikes up a mountain 1 mi/h slower than she hikes down the mountain. if it takes her 1.5 hours to hike up the mountain, and only 1 hour to hike down it, how fast does she movein each direction?
..
let speed down the mountain be x mph
up the mountain speed = x-1 mph.
..
time down the mountain = 1 hour
time up the mountain = 1.5 hours
..
Distance up & down is the same.
..
distance = speed * time
1*x= (x-1)*1.5
x=1.5x-1.5
1.5x-x = 1.5
0.5x = 1.5
x= 1.5/0.5
x = 3 mph speed down the mountain.
3-1 = 2mph speed up the mountain.